OpenStudy (anonymous):

Simplify by factoring Square Root of 90

OpenStudy (anonymous):

\[\sqrt{90}=\sqrt{9*10}=\sqrt{9}*\sqrt{10}=3*\sqrt{10}\]

OpenStudy (anonymous):

Thank you!! I don't get math at all!!

OpenStudy (anonymous):

welcome ^_|_^

OpenStudy (anonymous):

May I ask you another question?